OpenStudy (anonymous):

solve each equation. 1/b+1 +1/b-1 = 2/b^2-1 PLEASE HELP WILL GIVE MEDAL.

OpenStudy (anonymous):

\[\frac{ 1 }{ b+1 } + \frac{ 1 }{ b-1 } = \frac{ 2 }{ b^2-1 }\]

OpenStudy (superman499):

wat does be equal

DivineSolar (divinesolar):

Here you go ill provide a picture of what and how i solved your problem.

DivineSolar (divinesolar):

There is the work for your problem.

DivineSolar (divinesolar):

All explained in the PNG

DivineSolar (divinesolar):

How is that @imogen ?

OpenStudy (anonymous):

THANK YOU SO MUCH

DivineSolar (divinesolar):

Sure thing.

DivineSolar (divinesolar):

By the way you don't need to give me a medal but if you want you can, as i don't want to break the coc for partaking in medal parties im just here to help.